OpenStudy (anonymous):

Solve the equation: √x+4=7

OpenStudy (yttrium):

hey @charlieberzak , since I taught you some tips. Can you show me some solutions, you've got from this? :D

OpenStudy (anonymous):

\[\sqrt{x+4}=7\]

OpenStudy (anonymous):

thats how is should look like lol

OpenStudy (anonymous):

soo i should bring something to the other side?

OpenStudy (anonymous):

thats what u said last time...

OpenStudy (yttrium):

No because it is also a part of the exponent. You square first the whole equation before transposing. :)

OpenStudy (anonymous):

O_o

OpenStudy (anonymous):

just square both side and separate variable and numbers

OpenStudy (yttrium):

As what we've said, you need to cancel the sqrt sign right? Therefore, what should you do first?

OpenStudy (anonymous):

ummm find x? which i think would be 3?

OpenStudy (yttrium):

No, square 7 first. Before subtracting 4. :)

OpenStudy (anonymous):

rootx = 7-4 x square = 3 square , x = 9

OpenStudy (anonymous):

what do you mean by square?

OpenStudy (yttrium):

multiply 7 by itself.

OpenStudy (anonymous):

\[x ^{2}\]

OpenStudy (anonymous):

7*7 = 49, is that what you mean?

OpenStudy (anonymous):

\[\sqrt{x} = 7 - 4\]

OpenStudy (anonymous):

the answer is 45!

OpenStudy (anonymous):

\[\sqrt{x} = 3\]

OpenStudy (yttrium):

Yes.

OpenStudy (anonymous):

O.O

OpenStudy (yttrium):

Yes. You're right @charlieberzak , @JOELMATHEWS1234 maybe you misinterpreted the question.

OpenStudy (anonymous):

\[x = 9\]

OpenStudy (anonymous):

give the full question

OpenStudy (yttrium):

it's written in one of the comments.

OpenStudy (anonymous):

oh , okay

OpenStudy (anonymous):

\[\sqrt{x + 4} = 7\]

OpenStudy (anonymous):

\[x + 4 = 7^{2}\]

OpenStudy (anonymous):

\[x = 49 -4 =45\]